OpenStudy (anonymous):

please solve using quadratic formula: 2r^2=9r+81

OpenStudy (anonymous):

Write the quadratic formula here

OpenStudy (anonymous):

-b±√ ̅ b^2-4ac _______________ 2a

OpenStudy (anonymous):

apply the formula, and write what you got

OpenStudy (anonymous):

1) -9±√ ̅ 9^2-4*2*81 _______________ 2*2 2)-9±√ ̅ 12474 _______________ 4 but then i get a realy big decimal

OpenStudy (anonymous):

what am i doing wrong?

OpenStudy (anonymous):

here is your mistake.. Make the equation =0 not as it is now. This will give you the equation.... \[2r^2-9r-81=0\] from here do the formula. \[\frac{ 9\pm \sqrt{(-9)^2-4(2)(-81)}}{ 4} \] \[\frac{ 9\pm \sqrt{729} }{ 4 }\] \[=\frac{ 9\pm27 }{ 4}\]

OpenStudy (anonymous):

ooooohh duhh stupid mistake lol thank you so much!!

OpenStudy (anonymous):

no problem